OpenStudy (anonymous):

Find the probability of rolling a 4 or an even number on a number cube. Enter your answer in simplified fraction form; example: 3/5.

OpenStudy (anonymous):

Did you copy the question correctly?

OpenStudy (anonymous):

"or" means add the probabilities, but because 4 is an odd number, its probability is already included in even numbers

OpenStudy (anonymous):

but because 4 is an even** number

OpenStudy (anonymous):

P (4 Or an even no.)=\[P(4)\cup P( an even no.)=P \left( 4 \right)+P \left( an even no. \right)-P \left( 4 and an even no \right)\]

OpenStudy (anonymous):

My question exactly, probabilty of even # or roliing a 4 is how the instructor has phrased the ?

OpenStudy (anonymous):

but spelled better!

OpenStudy (anonymous):

Probability of getting 4=1/6 probability of getting an even no=3/6 probability of getting 4 and an even number=1/6 Required probability=1/6+3/6-1/6=3/6=1/2